OpenStudy (anonymous):

The degree measure of an angle is theta is between 0deg and 180 deg.If cosine of 0=1/2 what does 0 equal to?

OpenStudy (anonymous):

\[\cos(\theta)=\frac{ 1 }{ 2 }=>\theta=60\]

OpenStudy (jhannybean):

take 0 as theta. \[\large \cos(\theta)= \frac12\]\[\large \theta = \cos^{-1}(\frac12)\] what angle measurement of cosine is 1/2?

OpenStudy (anonymous):

sine

OpenStudy (jhannybean):



OpenStudy (luigi0210):

She meant what is the angle that creates this triangle